Can someone help me on this problem?
I need to solve and show proof on the following equation:
5x + 8 + 3x -x + 5 = 6x -3
Thanks,
Lisa

Answer by funmath(2873)

I'm not sure what is meant by "show proof", I'm going to solve it and tell you which property I used and then check it by substitution. I hope that's what you need.
5x+8+3x-x+5=6x-3
5x+3x-x+8+5=6x-3 Commutative Property of addition
(5x+3x-x)+(8+5)=6x-3 Associative property of addition
(5+3-1)x+(8+5)=6x-3 distributive propty
7x+13=6x-3 Simplification
-6x+7x+13=-6x+6x-3 Additive inverse
(-6+7)x+13=(-6+6)x-3 Distributive Property
x+13=-3 simplification
x+13-13=-3-13 additative inverse
x=-16
-----------------------------------------
Check:
5(-16)+8+3(-16)-(-16)+5=6(-16)-3
-80+8-48+16+5=-96-3
-99=-99
We're Correct!!!
Happy calculating!!!